# EIAD: Explainable Industrial Anomaly Detection Via Multi-Modal Large Language Models

### v1

@@ Abstract
Industrial Anomaly Detection (IAD) is critical to ensure product quality during manufacturing. Although existing zero-shot defect segmentation and detection methods have shown effectiveness, they cannot provide detailed descriptions of the defects. Furthermore, the application of large multi-modal models in IAD remains in its infancy, facing challenges in balancing question-answering (QA) performance and mask-based grounding capabilities, often owing to overfitting during the fine-tuning process. To address these challenges, we propose a novel approach that introduces a dedicated multi-modal defect localization module to decouple the dialog functionality from the core feature extraction. This decoupling is achieved through independent optimization objectives and tailored learning strategies. Additionally, we contribute to the first multi-modal industrial anomaly detection training dataset, named Defect Detection Question Answering (DDQA), encompassing a wide range of defect types and industrial scenarios. Unlike conventional datasets that rely on GPT-generated data, DDQA ensures authenticity and reliability and offers a robust foundation for model training. Experimental results demonstrate that our proposed method, Explainable Industrial Anomaly Detection Assistant (EIAD), achieves outstanding performance in defect detection and localization tasks. It not only significantly enhances accuracy but also improves interpretability. These advancements highlight the potential of EIAD for practical applications in industrial settings.
@@ I Introduction
Industrial anomaly detection plays a pivotal role in ensuring the reliability and safety of modern manufacturing systems. As industrial processes become increasingly complex and the diversity of potential defects grows, existing methods[1, 2, 3, 4] that primarily rely on predicting pixel-level anomaly scores are limited in their ability to define and describe anomalous regions precisely. Interpretative details related to domain knowledge such as anomaly categories, visual semantics, and potential causes and consequences are often overlooked. Recently, large multi-modal models have shown promising capabilities across various domains by integrating visual and language understanding to solve complex tasks. However, their application in industrial anomaly detection remains in its infancy, and several critical challenges remain to be addressed.
@@
Figure 1: Examples of four tasks from the DDQA-set
@@ II-B Large Visual language models(LVLMs)
Large Vision-Language Models (LVLMs) integrate visual modalities into large language models and have demonstrated exceptional perception and generalization capabilities across a wide range of vision tasks including image captioning, visual question answering, cross-modal retrieval and specific downstream tasks like camouflaged object detection[6, 7, 8, 9, 10, 11]. Recent studies have increasingly focused on enhancing the fine-grained grounding capabilities of LVLMs. For instance, methods such as LISA[12], GLaMM[13], GSVA[14], and SAM4MLLM[15] incorporate specific vocabularies into models and leverage decoding modules to generate mask grounding outputs. These approaches typically employ grounding supervision during fine-tuning to further refine the localization performance of LVLMs.
@@ III-A Construction of the Proposed DDQA-set
TABLE I: Statistical information of DDQA-set
@@
TABLE II: Comparison of dataset difficulty on DDQA and MMAD
@@ IV-B Implementation Details
On the DDQA-Set, we initially fine-tuned the EDDM module using LoRA (rank=8, alpha=32), such as MiniCPM-V2.6[21]. The model is trained on NVIDIA A100 80G GPUs by the SWIFT framework. Afterward, we fine-tune the MDLM module (based on LLavA-v1.6[7]) and Segment Anything Model[23] with LoRA (rank=16, alpha=16), training for 200 steps on the same hardware configuration, with a learning rate of 2e-3 and a batch size of 8.
@@ IV-C Comparison with LVLMS
To comprehensively evaluate the quality of defect detection, we conducted experiments using the MMAD dataset proposed in[22] to validate the effectiveness of our method. As shown in Table III, our approach demonstrates superior performance across all tasks in the MMAD dataset. Our EIAD is comparable to larger open-source models and most of the closed-source models.
@@ IV-D Comparison with Anomaly Detection Methods
TABLE IV: Zero-shot anomaly detection results on MVTec AD[25]. The best-performing method is in bold.
Methods Base Model MVTec AD
AUROC F1-max AP-seg
WinCLIP ViT-B/16+ 85.1 31.7 18.2
SAA+ GroundingDINO 72.2 37.8 28.8
AnoCLIP ViT-B/16+ 86.6 30.1 -
CLIP Surgery ViT-B/16+ 83.5 29.8 -
EIAD (ours) LLaVA (7B) 87.8 34.7 29.7
TABLE V: Zero-shot anomaly detection results on ViSA[26] dataset. The best-performing method is in bold.
Methods Base Model ViSA
AUROC F1-max AP-seg
WinCLIP ViT-B/16+ 79.6 14.8 5.4
SAA+ GroundingDINO 74.0 27.1 22.4
AnoCLIP ViT-B/16+ 83.7 13.5 -
CLIP Surgery ViT-B/16+ 85.0 15.2 -
EIAD (ours) LLaVA (7B) 91.9 18.7 14.0
To validate the generalization ability of our EIAD, we also use a 3D anomaly detection dataset, MVTec 3D-AD[27], to test the segmentation results of EIAD on its RGB images. We compared the our EIAD model with the baseline method WinCLIP. The results are presented in Table VI.
TABLE VI: Zero-shot anomaly detection results on MVTec 3D-AD[27].The best-performing method is in bold.
@@ IV-E Ablation Study
As shown in Table VII, the absence of the EDDM module led to a measurable decline in the quality of abnormal segmentation. This result demonstrates that the detailed defect descriptions generated by the EDDM module significantly enhance the grounding output.
TABLE VII: Ablation studies on the zero-shot anomaly segmentation task with and without EDDM module.

### v2

@@ Abstract
Industrial Anomaly Detection (IAD) is critical to ensure product quality during manufacturing. Although existing zero-shot defect segmentation and detection methods have shown effectiveness, they cannot provide detailed descriptions of the defects. Furthermore, the application of large multi-modal models in IAD remains in its infancy, facing challenges in balancing question-answering (QA) performance and mask-based grounding capabilities, often owing to overfitting during the fine-tuning process. To address these challenges, we propose a novel approach that introduces a dedicated multi-modal defect localization module to decouple the dialog functionality from the core feature extraction. This decoupling is achieved through independent optimization objectives and tailored learning strategies. Additionally, we contribute to the first multi-modal industrial anomaly detection training dataset, named Defect Detection Question Answering (DDQA), encompassing a wide range of defect types and industrial scenarios. Unlike conventional datasets that rely on GPT-generated data, DDQA ensures authenticity and reliability and offers a robust foundation for model training. Experimental results demonstrate that our proposed method, Explainable Industrial Anomaly Detection Assistant (EIAD), achieves outstanding performance in defect detection and localization tasks. It not only significantly enhances accuracy but also improves interpretability. These advancements highlight the potential of EIAD for practical applications in industrial settings. Code and datasets will be available at https://github.com/Solunny/EIAD.
@@ I Introduction
Industrial anomaly detection plays a pivotal role in ensuring the reliability and safety of modern manufacturing systems. As industrial processes become increasingly complex and the diversity of potential defects grows, existing methods[1, 2] that primarily rely on predicting pixel-level anomaly scores are limited in their ability to define and describe anomalous regions precisely. Interpretative details related to domain knowledge such as anomaly categories, visual semantics, and potential causes and consequences are often overlooked. Recently, large multi-modal models have shown promising capabilities across various domains by integrating visual and language understanding to solve complex tasks. However, their application in industrial anomaly detection remains in its infancy, and several critical challenges remain to be addressed.
@@
Figure 1: Examples of four tasks from the DDQA-set.
@@ II-B Large Visual language models (LVLMs)
Large Vision-Language Models (LVLMs) integrate visual modalities into large language models and have demonstrated exceptional perception and generalization capabilities across a wide range of vision tasks including image captioning, visual question answering and cross-modal retrieval[4, 5]. Recent studies have increasingly focused on enhancing the fine-grained grounding capabilities of LVLMs. For instance, methods such as LISA[6], GLaMM[7], and SAM4MLLM[8] incorporate specific vocabularies into models and leverage decoding modules to generate mask grounding outputs. These approaches typically employ grounding supervision during fine-tuning to refine the localization performance of LVLMs further.
@@ III-A Construction of the Proposed DDQA-set
TABLE I: Statistical information of DDQA-set.
@@
TABLE II: Comparison of dataset difficulty on DDQA and MMAD.
@@ IV-B Implementation Details
On the DDQA-Set, we initially fine-tuned the EDDM module using LoRA (rank=8, alpha=32), such as MiniCPM-V2.6[14]. Afterward, we fine-tune the MDLM module (based on LLavA-v1.6[4]) and Segment Anything Model[16] with LoRA (rank=16, alpha=16), training for 200 steps on the same hardware configuration, with a learning rate of 2e-3 and a batch size of 8.
@@ IV-C Comparison with LVLMs
To comprehensively evaluate the quality of defect detection, we conducted experiments using the MMAD dataset proposed in[15] to validate the effectiveness of our method. As shown in table III, our approach demonstrates superior performance across all tasks in the MMAD dataset and is comparable to larger open-source models and most of the closed-source models.
@@ IV-D Comparison with Anomaly Detection Methods
TABLE IV: Zero-shot anomaly detection results on MVTec AD[18] and ViSA[19]. The best-performing method is in bold and the second best score is underlined.
Methods MVTec AD ViSA
AUROC F1-max AUROC F1-max
WinCLIP 85.1 31.7 79.6 14.8
SAA+ 72.2 37.8 74.0 27.1
AnoCLIP 86.6 30.1 83.7 13.5
CLIP Surgery 83.5 29.8 85.0 15.2
EIAD (ours) 87.8 34.7 91.9 18.7
To validate the generalization ability of our EIAD, we also use a 3D anomaly detection dataset, MVTec 3D-AD[20], to test the segmentation results of EIAD on its RGB images. We compared the our EIAD model with the baseline method WinCLIP. The results are presented in table V.
TABLE V: Zero-shot anomaly detection results on MVTec 3D-AD[20].The best-performing method is in bold.
@@ IV-E Ablation Study
As shown in table VI, the absence of the EDDM module led to a measurable decline in the quality of abnormal segmentation. This result demonstrates that the detailed defect descriptions generated by the EDDM module significantly enhance the grounding output.
TABLE VI: Ablation studies on the zero-shot anomaly segmentation task with and without EDDM module.